# Pillar-Voxel Fusion Network for 3D Object Detection in Airborne Hyperspectral Point Clouds

### v1

@@ 1 Introduction
Hyperspectral Point Clouds (HPCs) acquired from airborne platforms integrate the advantages of LiDAR point clouds and hyperspectral imaging, providing rich spectral and spatial information[1, 2, 3, 4]. By integrating precise 3D measurements with detailed spectral signatures, HPCs enable accurate inversion of both material properties and spatial structures of surface objects. This unique combination of capabilities demonstrates significant potential across multiple applications, including 3D urban planning, precision agriculture, forest management, and camouflage detection.[5, 6, 7, 8, 9, 10, 11].
@@
Inspired by these challenges, we propose a 3D object detection network for airborne HPCs, named PiV-AHPC. We design an intuitive and effective pillar-voxel dual-branch encoder to extract complex heterogeneous features and overcome spectral distortion. Due to the advantage of the pillar structure’s vertical receptive field and its natural adaptability to the aerial view, this branch extracts spectral features for classification and determines vertical structure features as spectral confidence criteria. Simultaneously, since the voxel structure preserves the spatial structure details, this branch captures fine-grained three-dimensional spatial features for precise target localization. We implement a multi-level feature fusion mechanism that contains two modules to mitigate geometric distortion and enhance branch correlation. Specifically, intermediate-level branch correlations are established by the sparse fusion module to provide more information for subsequent bounding box refinement. During the output of the BEV feature map, the patch-wise fusion module achieves an organic integration of neighboring features through adaptive selection and alignment. This well-designed architecture ensures the synergy of the encoders and fully mines the rich information in airborne HPCs.
@@ 2.1 Overall architecture
First, we reduce the dimensionality of the HPCs by principal component analysis (PCA) and input the pillar branch after pillarization. Simultaneously, we split point clouds containing positional information from the HPCs and input the voxel branch after voxelization. During the feature encoding process in both branches, the middle-layer features are dynamically adjusted and integrated by a sparse feature fusion module. At the end of the two encoders, a patch-wise feature fusion module adaptively aligns and selects dual-branch neighborhood feature channels, obtaining a BEV representation of the scene information. Finally, this representation is fed into a detection head with the same structure as the Voxel-RCNN to produce the prediction results. The overall architecture is illustrated in Figure 3.
@@ 2.2 Pillar-Voxel Dual-Branch Encoder
Pillars can be viewed as BEV-projected representations, naturally adapting to airborne hyperspectral imagery. In addition, the occlusion of obstacles in the vertical direction is the primary cause of HPCs spectral distortion. The sufficient receptive field of pillars in the vertical direction facilitates the extraction of vertical structural features, which serve as the discrimination basis for object occlusion and enable the algorithm to adaptively utilize spectral features. Finally, the excellent computational efficiency of pillars ensures the algorithm’s running speed when handling redundant HPCs.
@@
Based on this analysis,integrating the structural advantages of both pillars and voxels is an intuitive strategy for extracting airborne HPCs features. Therefore, we design a pillar-voxel dual-branch encoder, where the pillar branch extracts spectral and vertical structural features, while the voxel branch captures 3D spatial structural features, achieving complementary advantages.
@@
In the pillar branch, a lightweight PointNet encoder first processes point clouds divided into identical pillars, incorporating features such as absolute coordinates, normalized coordinates, center points, elevation range, and spectral data to obtain the initial features ${{F}_{p}}\in{{\mathbb{R}}^{{{N}_{p}}\times{{C}_{p}}}}$. This branch consists of four stages: the first stage uses two layers of submanifold sparse convolution [52] for preprocessing, while the remaining stages combine one layer of 2D sparse convolution with two layers of submanifold sparse convolution to form a 2D sparse convolution block (SCB), as shown in Figure 3. Efficient sparse convolution operators downsample the scene, reducing computational load by focusing on relevant features and ignoring redundant spatial data. Submanifold convolution maintains data sparsity and structural integrity by restricting output to existing non-empty elements. Finally, the 8$\times$×downsampled sparse features $F_{p}^{4}$ are transformed into dense pillar features $F_{p}^{D}\in{{\mathbb{R}}^{H\times W\times C}}$ based on indices $I_{p}^{4}$.
In the voxel branch, the mean coordinates of the point clouds within the same voxel are calculated to obtain the initial features ${{F}_{v}}\in{{\mathbb{R}}^{{{N}_{v}}\times 3}}$. Similar to the pillar branch, a four-stage 3D encoder is employed in the first half of the voxel branch. At the end of each stage, focal convolution [38] is introduced, which adaptively changes kernel shapes during training to retain valuable foreground information, forming focal sparse convolution blocks (FSB). To ensure consistent resolution and similar levels of abstraction in the horizontal direction for the intermediate layer features extracted by the dual branches, the corresponding SCB and FSB maintain the same horizontal convolution kernel size, stride, padding, and output feature dimensions. This facilitates the subsequent element-wise alignment and sparse fusion of features. Ultimately, 8$\times$×downsampled sparse voxels $F_{v}^{4}\in{{\mathbb{R}}^{N_{v}^{4}\times C_{v}^{4}}}$ are obtained.
@@
To address these issues, we introduce a learnable parameter $z\in{{\mathbb{R}}^{D\times 1}}$, where D represents the number of vertical dimensions in the scene, is introduced. The weight sparsity $W\in{{\mathbb{R}}^{D\times 1}}$assigned to different elevations is obtained as follows:
@@
is the set of voxels at the same BEV position, and Unique represents the de-duplication operator. Finally, $F_{v}^{b}$ is transformed into dense voxel features $F_{v}^{D}\in{{\mathbb{R}}^{H\times W\times C}}$.
@@ 2.3 Multi-Level Feature Fusion Mechanism
Compared to a vanilla dual-branch encoder without feature exchange during extraction, sparse feature fusion in intermediate layers establishes information exchange between branches, enabling the network to dynamically adjust feature correlations. Furthermore, the spectral and vertical structure information extracted by the pillar branch enriches the intermediate voxel features, which provides spectral confidence criterion and a more effective classification basis for the subsequent RoI pooling layer.
@@ 3.2 Experimental Setup
2)Implementation Details: Our algorithm was implemented using the OpenPCDet codebase based on the PyTorch framework. For both datasets, we employed the DistributeDataParallel strategy to train the network on four TitanXP GPUs, with a batch size of 2 per GPU. We used the Adam optimizer with a learning rate and weight decay of 0.01, training for a total of 30 epochs. The input voxel size was set to (0.1m, 0.1m, 0.1m), the input pillar size was set to (0.1m, 0.1m, 16m), and the detection range was configured as (-25.6 25.6m, -25.6 25.6m, 0 16m).
@@ 3.3 Comparative Analysis with the State-of-the-Art
3.4 Comparative Analysis of Different Occlusion Levels
TABLE 6,7 display the object detection accuracy of PiV-AHPC under different levels of occlusion in two datasets. The HIT Campus Dataset does not include part occlusion tents, as well as occlusion-free tents and boxes. According to the table content, the algorithm’s detection accuracy shows a low correlation with the occlusion level, suggesting that occlusion has little impact on PiV-AHPC’s performance. Specifically, in the HIT Campus Dataset, compared to part occlusion and occlusion-free scenes, the algorithm in high occlusion scene for vehicle objects achieves improvements of 1.75% and 5.23% in $\text{A}{{\text{P}}_{\text{BEV}}}$, and 3.76% and 11.82% in $\text{A}{{\text{P}}_{\text{3D}}}$. For part occlusion box objects, improvements in $\text{A}{{\text{P}}_{\text{BEV}}}$ and $\text{A}{{\text{P}}_{\text{3D}}}$ compared to high occlusion objects are 1.74% and 2.14%, respectively. The algorithm achieves the best detection accuracy for high occlusion objects in the Houston2018 dataset, with $\text{A}{{\text{P}}_{\text{3D}}}$ surpassing the other two types of objects by 2.56% and 1.18%, respectively. Spectral assignments of low-level objects in HPCs are often incorrect when occlusion occurs. However, PiV-AHPC demonstrates promising performance in high-occlusion scenes across two datasets. On the one hand, these results highlight the advantages of LiDAR, including its ability to penetrate occlusions and achieve 3D imaging. On the other hand, they demonstrate the effectiveness of PiV-AHPC’s pillar branch, which has a broad vertical receptive field. This branch selectively utilizes spectral information based on the distribution of specific objects, suppressing interference from incorrect spectral values.
Method Difficulty Vehicle Tent Box mAP
BEV 3D AOS BEV 3D AOS BEV 3D AOS BEV 3D AOS
Mod 52.88 22.02 52.66 73.36 7.89 35.14 24.3 2.67 10.8 50.18 10.86 32.87
Hard 40.09 16 39.76 40.22 3.92 19.2 17.76 1.84 7.78 32.69 7.25 22.24
Mod 43.26 9.35 41.6 72.65 63.49 67.75 23.26 7.34 18.95 46.85 26.77 43
Hard 34.8 7.03 32.93 39.94 36.36 37.16 20.13 6.5 16.39 30.97 15.28 28
Hard 41.74 18.68 41.41 37.56 9.93 26.05 30.24 4.15 19.83 36.52 10.92 29.1
Hard 43.21 14 42.82 47.83 5.84 24.5 11.66 1.09 5.44 34.24 6.98 24.25
Method $\text{A}{{\text{P}}_{\text{BEV}}}$ $\text{A}{{\text{P}}_{\text{3D}}}$ AOS
Easy Mod Hard 0.5:0.95 Easy Mod Hard 0.5:0.95 Easy Mod Hard 0.5:0.95
PillarNet 18.89 25.76 26.74 15.19 0.34 0.32 0.27 3.55 9.51 12.71 13.43 7.60
Method $\text{mA}{{\text{P}}_{\text{BEV}}}$ $\text{mA}{{\text{P}}_{\text{3D}}}$ AOS
Easy Mod Hard Easy Mod Hard Easy Mod Hard
PartA2 70.88 68.32 65.26 46.22 44.45 43.19 51.37 50.31 49.07
CenterPoint 78.7 76.37 73.21 25.29 23.89 21.96 49.95 47.63 44.47
PillarNet 18.31 17.83 17.3 6.32 6.16 6.13 12.37 12.02 11.67
PV-RCNN++ 63.09 60.62 57.32 52.64 50.43 48.18 54.23 51.76 48.44
Voxel-RCNN 77.23 74.83 71.7 57 55.56 53.77 62.85 61.12 59.12
TransFusion-L 56.5 56.17 56.11 5.75 5.74 5.74 28.34 28.18 28.15
VoxelNext 78.93 74.84 71.56 41.47 38.63 36.36 53.52 49.75 46.48
PiV-AHPC 80.12 78.44 75.19 57.4 55.46 54.19 70.57 68.85 65.77
@@ 3.6 Ablation Study
We conducted ablation studies on several key modules of PiV-AHPC on the HIT Campus Dataset to evaluate their effectiveness.
1)Effectiveness of pillar-voxel Encoder: As the starting point of the network, the pillar-voxel dual-branch encoder is the foundation of the proposed method in this work. To test its effectiveness, we employed Voxel-RCNN as a baseline for single-branch networks and sequentially added the pillar branch, the sparse feature fusion module. TABLE 8 presents the mAP scores of each algorithm. The results indicate that our dual-branch encoder significantly enhances the detection capabilities of the original algorithm under three difficulty conditions. Taking the ‘Hard’ level of BEV, 3D, and angle metrics as examples: the initial precision of the vanilla network was 42.24%, 21.01%, and 38.91%, respectively. After introducing the pillar branch, these metrics improved by 5.21%, 1.82%, and 5.52%, respectively. Further addition of the sparse feature fusion module to facilitate information exchange between branches resulted in a continued improvement in detection accuracy by 2.74%, 4.52%, and 2.46%. These ablation experiment results demonstrate the significant advantage of our proposed dual-branch encoder in capturing the complex spatial and spectral relationships between HPCs data. And this advantage will be further amplified by establishing inter-branch correlations.
2)Effectiveness of Key Modules: In addition to the dual-branch encoder, we designed the patch-wise feature fusion (P) module as well as the multi-scale feature aggregation module with additional downsampling layers (A) and weighted elevation compression (W). We conducted ablation studies on these key modules, using the algorithm’s detection accuracy at the ‘Hard’ level as the evaluation criterion, as shown in TABLE 9. It can be observed that the absence of any module leads to a decrease in detection capability to varying degrees. Specifically, removing the ‘P’ resulted in a 2.2% decrease in $\text{mA}{{\text{P}}_{\text{3D}}}$, while discarding the ’W’ module led to a 3.34% decrease. Eliminating the ’A’ module resulted in a significant accuracy drop to 29.34%. These experiments illustrate the significant role of the proposed key modules. The multi-scale feature aggregation module, with its diverse 3D receptive fields and selective elevation compression, significantly enhanced model performance. Although the patch-wise feature fusion module had limited impact on accuracy, its Transformer architecture demonstrated potential for application on large-scale datasets, offering possibilities for further enhancing model performance.

### v2

@@ 1 Introduction
Hyperspectral Point Clouds (HPCs) acquired from airborne platforms integrate the advantages of LiDAR point clouds and hyperspectral imaging, providing rich spectral and spatial information[1, 2, 3, 4]. By integrating precise 3D measurements with detailed spectral signatures, HPCs enable accurate inversion of both material properties and spatial structures of surface objects. This unique combination of capabilities demonstrates significant potential across multiple applications, including 3D urban planning, precision agriculture, forest management, and camouflage detection[5, 6, 7, 8, 9, 10].
@@
Inspired by these challenges, we propose a Pillar-Voxel Fusion Network for 3D Object Detection in Airborne Hyperspectral Point Clouds, named PiV-AHPC. We design an intuitive and effective pillar-voxel dual-branch encoder to extract complex heterogeneous features and overcome spectral distortion. Due to the advantage of the pillar structure’s vertical receptive field and its natural adaptability to the BEV, this branch extracts spectral features for classification and determines vertical structure features as spectral confidence criteria. Simultaneously, since the voxel structure preserves the spatial structure details, this branch captures fine-grained three-dimensional spatial features for precise target localization. The pillar branch extracts vertical structural features that focus on capturing point cloud distribution along the z-axis, while the voxel branch extracts 3D spatial features that emphasize capturing fine 3D geometric details. We implement a multi-level feature fusion mechanism that contains two modules to mitigate geometric distortion and enhance branch correlation. Specifically, intermediate-level branch correlations are established by the sparse fusion module to provide more information for subsequent bounding box refinement. During the output of the BEV feature map, the patch-wise fusion module achieves an organic integration of neighboring features through adaptive selection and alignment. This well-designed architecture ensures the synergy of the encoders and fully mines the rich information in airborne HPCs.
@@ 2.1 Overall architecture
First, we reduce the spectral dimensions of HPCs through Principal Component Analysis (PCA, default reduction to 21 components), and input the pillar branch after pillarization. This PCA operation can eliminate redundant information and reduce computational resource consumption. Simultaneously, we split point clouds containing positional information from the HPCs and input the voxel branch after voxelization. During the feature encoding process in both branches, the middle-layer features are dynamically adjusted and integrated by a sparse feature fusion module. At the end of the two encoders, a patch-wise feature fusion module adaptively aligns and selects dual-branch neighborhood feature channels, obtaining a BEV representation of the scene information. Finally, this representation is fed into a detection head with the same structure as the Voxel-RCNN to produce the prediction results. The overall architecture is illustrated in Figure 3.
@@ 2.2 Pillar-Voxel Dual-Branch Encoder
Pillars can be viewed as BEV-projected representations, naturally adapting to airborne hyperspectral imagery. In addition, due to the inability of passive hyperspectral imaging to penetrate obstacles (such as tree canopies), occluded objects and edges inevitably exhibit spectral distortion. Pillars possess sufficient receptive field in the vertical direction, allowing them to capture structural information throughout the entire vertical space, thereby identifying spectral distortion areas where occlusion exists. Finally, the excellent computational efficiency of pillars ensures the algorithm’s running speed when handling redundant HPCs.
@@
Based on this analysis,integrating the structural advantages of both pillars and voxels is an intuitive strategy for extracting airborne HPCs features. Therefore, we design a pillar-voxel dual-branch encoder, where the pillar branch extracts spectral and vertical structural features, while the voxel branch captures 3D spatial structural features, achieving complementary advantages. The algorithm uses the vertical structural features extracted by the pillar branch as criteria for spectral reliability, thereby evaluating the reliability of spectral data in different regions. When an area with spectral distortion is identified, the network adaptively reduces its dependence on the spectral features of that region, instead relying more on the fine-grained spatial structural features extracted by the voxel branch. This adaptive mechanism effectively mitigates the interference of spectral distortion on detection performance.
@@
In the pillar branch, a lightweight PointNet encoder first processes point clouds divided into identical pillars, incorporating features such as absolute coordinates, normalized coordinates, center points, elevation range, and spectral data to obtain the initial features ${{F}_{p}}\in{{\mathbb{R}}^{{{N}_{p}}\times{{C}_{p}}}}$, where $C_{p}$ represents the number of feature channels in the Pillar branch. This branch consists of four stages: the first stage uses two layers of submanifold sparse convolution [51] for preprocessing, while the remaining stages combine one layer of 2D sparse convolution with two layers of submanifold sparse convolution to form a 2D sparse convolution block (SCB), as shown in Figure 3. Efficient sparse convolution operators downsample the scene, reducing computational load by focusing on relevant features and ignoring redundant spatial data. Submanifold convolution maintains data sparsity and structural integrity by restricting output to existing non-empty elements. Finally, the 8$\times$×downsampled sparse features $F_{p}^{4}$ are transformed into dense pillar features $F_{p}^{D}\in{{\mathbb{R}}^{H\times W\times C}}$ based on indices $I_{p}^{4}$.
In the voxel branch, the mean coordinates of the point clouds within the same voxel are calculated to obtain the initial features ${{F}_{v}}\in{{\mathbb{R}}^{{{N}_{v}}\times 3}}$. Similar to the pillar branch, a four-stage 3D encoder is employed in the first half of the voxel branch. At the end of each stage, focal convolution [37] is introduced, which adaptively changes kernel shapes during training to retain valuable foreground information, forming focal sparse convolution blocks (FSB). To ensure consistent resolution and similar levels of abstraction in the horizontal direction for the intermediate layer features extracted by the dual branches, the corresponding SCB and FSB maintain the same horizontal convolution kernel size, stride, padding, and output feature dimensions. This facilitates the subsequent element-wise alignment and sparse fusion of features. Ultimately, 8$\times$×downsampled sparse voxels $F_{v}^{4}\in{{\mathbb{R}}^{N_{v}^{4}\times C_{v}^{4}}}$ are obtained, where $C_{v}^{4}$ represents the number of feature channels in the fourth layer of the Voxel branch.
@@
To address these issues, we introduce a learnable parameter $z\in{{\mathbb{R}}^{D\times 1}}$, where $D$ represents the number of dimensions in the vertical direction of the voxel features. The sparse weights $W\in{{\mathbb{R}}^{D\times 1}}$ for different heights are obtained through Softmax normalization of the learnable parameter $z$, with the meaning of importance weights of voxels at different vertical heights for detection scene:
@@
represents the set of all vertical voxel features at the same horizontal coordinates $(x,y)$ in the BEV, Unique represents the de-duplication operator, and $N_{b}$ represents the number of unique BEV positions retained after height compression, which is the total number of deduplicated horizontal coordinates $(x,y)$. Finally, $F_{v}^{b}$ is transformed into dense voxel features $F_{v}^{D}\in{{\mathbb{R}}^{H\times W\times C}}$.
@@ 2.3 Multi-Level Feature Fusion Mechanism
Compared to a vanilla dual-branch encoder without feature exchange during extraction, sparse feature fusion in intermediate layers establishes information exchange between branches, enabling the network to dynamically adjust feature correlations. Furthermore, the spectral and vertical structure information extracted by the pillar branch enriches the intermediate voxel features, which provides spectral confidence criterion and a more effective classification basis for the subsequent refinement network.
@@ 3.2 Experimental Setup
2)Implementation Details: Our algorithm was implemented using the OpenPCDet[61] codebase based on the PyTorch framework. For both datasets, we employed the Distributed Data Parallel strategy (4×TitanXP GPUs, batch size of 2 per GPU), AdamW optimizer (learning rate 0.005, weight decay 0.01), and OneCycle scheduling strategy (peak learning rate 0.05, momentum range [0.85, 0.95]), training for a total of 30 epochs. The loss functions are structured in two network stages: the region proposal network employs Focal classification loss, Smooth L1 regression loss, and direction classification loss, while the refinement network utilizes cross-entropy classification loss, Smooth L1 regression loss, and 3D corner regression loss. All hyperparameters for these loss components follow the OpenPCdet settings. The input voxel size was set to (0.1m, 0.1m, 0.1m), the input pillar size was set to (0.1m, 0.1m, 16m), and the detection range was configured as (-25.6 25.6m, -25.6 25.6m, 0 16m).
@@ 3.3 Comparative Analysis with the State-of-the-Art
Method Difficulty Vehicle Tent Box mAP
BEV 3D AOS BEV 3D AOS BEV 3D AOS BEV 3D AOS
Mod 52.88 22.02 52.66 73.36 7.89 35.14 24.3 2.67 10.8 50.18 10.86 32.87
Hard 40.09 16 39.76 40.182 3.92 19.2 17.76 1.84 7.78 32.69 7.25 22.24
Mod 43.26 9.35 41.6 72.65 63.49 67.75 23.26 7.34 18.95 46.85 26.77 43
Hard 34.8 7.03 32.93 39.94 36.36 37.16 20.13 6.5 16.39 30.97 15.28 28
Hard 41.74 18.68 41.41 37.56 9.93 26.05 30.184 4.15 19.83 36.52 10.92 29.1
Hard 43.21 14 42.82 47.83 5.84 24.5 11.66 1.09 5.44 34.24 6.98 24.25
@@
Method $\text{A}{{\text{P}}_{\text{BEV}}}$ $\text{A}{{\text{P}}_{\text{3D}}}$ AOS
Easy Mod Hard 0.5:0.95 Easy Mod Hard 0.5:0.95 Easy Mod Hard 0.5:0.95
PillarNet 18.89 25.76 26.74 15.19 0.34 0.32 0.187 3.55 9.51 12.71 13.43 7.60
3.4 Comparative Analysis of Different Occlusion Levels
TABLE 6,7 display the object detection accuracy of PiV-AHPC under different levels of occlusion in two datasets. The HIT Campus Dataset does not include part occlusion tents, as well as occlusion-free tents and boxes. According to the table content, the algorithm’s detection accuracy shows a low correlation with the occlusion level, suggesting that occlusion has little impact on PiV-AHPC’s performance. Specifically, in the HIT Campus Dataset, compared to part occlusion and occlusion-free scenes, the algorithm in high occlusion scene for vehicle objects achieves improvements of 1.75% and 5.23% in $\text{A}{{\text{P}}_{\text{BEV}}}$, and 3.76% and 11.82% in $\text{A}{{\text{P}}_{\text{3D}}}$. For part occlusion box objects, improvements in $\text{A}{{\text{P}}_{\text{BEV}}}$ and $\text{A}{{\text{P}}_{\text{3D}}}$ compared to high occlusion objects are 1.74% and 2.14%, respectively. The algorithm achieves the best detection accuracy for high occlusion objects in the Houston2018 dataset, with $\text{A}{{\text{P}}_{\text{3D}}}$ surpassing the other two types of objects by 2.56% and 1.18%, respectively. Spectral assignments of low-level objects in HPCs are often incorrect when occlusion occurs. However, PiV-AHPC demonstrates promising performance in high-occlusion scenes across two datasets. On the one hand, these results highlight the advantages of LiDAR, including its ability to penetrate occlusions and achieve 3D imaging. On the other hand, they demonstrate the effectiveness of PiV-AHPC’s pillar branch, which has a sufficient vertical receptive field. This branch selectively utilizes spectral information based on the distribution of specific objects, suppressing interference from incorrect spectral values.
Method $\text{mA}{{\text{P}}_{\text{BEV}}}$ $\text{mA}{{\text{P}}_{\text{3D}}}$ AOS
Easy Mod Hard Easy Mod Hard Easy Mod Hard
PartA2 70.88 68.32 65.26 46.22 44.45 43.19 51.37 50.31 49.07
CenterPoint 78.7 76.37 73.21 25.29 23.89 21.96 49.95 47.63 44.47
PillarNet 18.31 17.83 17.3 6.32 6.16 6.13 12.37 12.02 11.67
PV-RCNN++ 63.09 60.62 57.32 52.64 50.43 48.18 54.23 51.76 48.44
Voxel-RCNN 77.23 74.83 71.7 57 55.56 53.77 62.85 61.12 59.12
TransFusion-L 56.5 56.17 56.11 5.75 5.74 5.74 28.34 28.18 28.15
VoxelNext 78.93 74.84 71.56 41.47 38.63 36.36 53.52 49.75 46.48
PiV-AHPC 80.12 78.44 75.19 57.4 55.46 54.19 70.57 68.85 65.77
@@ 3.5 Analysis of algorithm transferability
Figure 13: Performance of PiV-AHPC and cumulative explained variance under different numbers of principal components. For the HIT Campus dataset, we compare the results using 10, 21, 105, 189 principal components, and the original data. For the Houston2018 dataset, we compare 3, 12, 21, 34 principal components, and the original data.
@@ 3.6 Ablation Study
1)Principal Component Selection: We conducted ablation experiments on both HIT Campus and Houston2018 datasets, comparing detection performance under different numbers of principal components and their cumulative explained variance (reflecting information retention), as shown in Figure 13. For the HIT Campus dataset, when reduced to 21 principal components, the model achieved optimal detection performance under ’Hard’ level, with $\text{mA}{{\text{P}}_{\text{3D}}}$ reaching 35.39%, significantly outperforming the original data’s 32.26% and 10 principal components’ 31.8%. This indicates that excessive dimensions may contain redundant information, while over-reduction can lose fine-grained information, leading to decreased detection accuracy. 21 principal components both retain critical information (cumulative explained variance of 99.97%) and reduce computational resource consumption. For the Houston2018 dataset, data with 21 principal components performed best on the $\text{A}{{\text{P}}_{\text{BEV}}}$ metric, although its $\text{A}{{\text{P}}_{\text{3D}}}$ was slightly lower than those with 12 and 34 dimensions. However, considering consistency with the HIT Campus dataset configuration, we chose 21 dimensions as the unified setting. Appropriate PCA dimensionality reduction not only improves computational efficiency but also enhances detection performance by reducing redundant information, achieving a balance between detection accuracy and computational cost.
2)Effectiveness of pillar-voxel Encoder: As the starting point of the network, the pillar-voxel dual-branch encoder is the foundation of the proposed method in this work. To test its effectiveness, we employed Voxel-RCNN as a baseline for single-branch networks and sequentially added the pillar branch, the sparse feature fusion module. TABLE 8 presents the mAP scores of each algorithm. The results indicate that our dual-branch encoder significantly enhances the detection capabilities of the original algorithm under three difficulty conditions. Taking the ‘Hard’ level of BEV, 3D, and angle metrics as examples: the initial precision of the vanilla network was 42.24%, 21.01%, and 38.91%, respectively. After introducing the pillar branch, these metrics improved by 5.21%, 1.82%, and 5.52%, respectively. Further addition of the sparse feature fusion module to facilitate information exchange between branches resulted in a continued improvement in detection accuracy by 2.74%, 4.52%, and 2.46%. These ablation experiment results demonstrate the significant advantage of our proposed dual-branch encoder in capturing the complex spatial and spectral relationships between HPCs data. And this advantage will be further amplified by establishing inter-branch correlations.
@@
3)Effectiveness of Key Modules: In addition to the dual-branch encoder, we designed the patch-wise feature fusion (P) module as well as the multi-scale feature aggregation module with additional downsampling layers (A) and weighted elevation compression (W). We conducted ablation studies on these key modules, using the algorithm’s detection accuracy at the ‘Hard’ level as the evaluation criterion, as shown in TABLE 9. It can be observed that the absence of any module leads to a decrease in detection capability to varying degrees. Specifically, removing the ‘P’ resulted in a 2.2% decrease in $\text{mA}{{\text{P}}_{\text{3D}}}$, while discarding the ’W’ module led to a 3.34% decrease. Eliminating the ’A’ module resulted in a significant accuracy drop to 29.34%. These experiments illustrate the significant role of the proposed key modules. The multi-scale feature aggregation module, with its diverse 3D receptive fields and selective elevation compression, significantly enhanced model performance. Although the patch-wise feature fusion module had limited impact on accuracy, its Transformer architecture demonstrated potential for application on large-scale datasets, offering possibilities for further enhancing model performance.